Grade 12th passAlgebra

Find the amount to be paid at the end of 3 years in each case:
Principal = ₹ 7,500 at 5% p.a.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Given: - Principal (P)= ₹7500, Rate (R) = 5% p.a. and Time (T) = 3years.
If interest is calculated uniformly on the original principal throughout the loan period, it is called Simple interest (SI).
SI = (P × R × T)/100
= (7500 × 5 × 3)/ 100
= (75 × 5 × 3)/ 1
= ₹1125
Amount = (principal + SI)
= (7500 + 1125)
= ₹8625

Thanks